A coffee shop mixes Brazilian coffee worth $5 per kilogram with Turkish coffee worth $8 per kilogram. The mixture is to sell for $7 per kilogram. How much of the Turkish coffee should be used to make 300 kg of the mixture?
Answer by Alan3354(69443)

A coffee shop mixes Brazilian coffee worth $5 per kilogram with Turkish coffee worth $8 per kilogram. The mixture is to sell for $7 per kilogram. How much of the Turkish coffee should be used to make 300 kg of the mixture?
--------------
B = amount of Brazilian coffee
5B + 8(300-B) = 7*300
5B + 2400 - 8B = 2100
B = 100 kg of Brazilian
--> 200 kg of Turkish